OpenStudy (anonymous):

Solve: (4x - 5)(3x + 2) =0 {5, -2} {5/4, - 2/3} {-5/4, 2/3} no solution

OpenStudy (anonymous):

4x-5=0 or 3x+2=0 x=5/4 or -2/3

OpenStudy (anonymous):

(4X-5)(3X+2) =0 It is given that both factors are equal to zero so, we can take as 4x-5=0 3x+2= 0 4x=5 3x=-2 x= -2/3 x= 5/4 Here we'll get two values of x = 5/4 ,-2/3

OpenStudy (anonymous):

(4X-5)(3X+2) =0 It is given that both factors are equal to zero so, we can take as 4x-5=0 3x+2= 0 4x=5 3x=-2 x= -2/3 x= 5/4 Here we'll get two values of x = 5/4 ,-2/3

OpenStudy (anonymous):

direct answers is crime here

OpenStudy (callisto):

@Rohangrr Sorry but I have to say it. you're right that giving answer without workings is not good here. But you should also know that copying others answers and posting it directly is unethical. I mean no offense to you, just reminding.